OpenStudy (anonymous):

suppose f(x)=4x-2and g(x)=-2x+1 find the value of f(5)/g(-3) options A)1and 5/9, B) 2 and 4/7, C)-2, and D)2

OpenStudy (anonymous):

f(5) means evaluating f when you plug in 5 for x: 4(5)-2=20-2=18 Similarly, g(-3): -2(-3)+1=6+1=7 So, f(5)/g(-3)=18/7 or 2 and 4/7, B.